OpenStudy (kewlgeek555):

Which equation has a solution of n = -1

OpenStudy (kewlgeek555):

Choices will be shown below:

OpenStudy (anonymous):

n+1=0 is the equation

OpenStudy (kewlgeek555):

a. \[\frac{ 1 }{ 2 }n - 6 = -7\] b. \[-6n + 8 = -2\] c. \[2(n - 8) = 18\] d. \[2n + 12 - 3n = 15\]

OpenStudy (kewlgeek555):

I picked d, but it wasn't right.

OpenStudy (kewlgeek555):

I true to solve all the problems, but it came out that d was right, but unfortunately it wasn't.

OpenStudy (anonymous):

are you sure they are asking for -1

OpenStudy (kewlgeek555):

Yes, it says, "Which equation has a solution for n = -1"

OpenStudy (anonymous):

okay so you solve all of them right what did you get

OpenStudy (kewlgeek555):

I got d, and so I submitted the test and it said I was wrong. Right now I am starting to solve them again while you guys see what you come up with.

OpenStudy (anonymous):

i think 'c' is relevant to that is lessthan or equal to

OpenStudy (kewlgeek555):

I solved a. It equals -0.5.

OpenStudy (anonymous):

that is the closet but it still don't equal one unless it is asking for less than or equal to

OpenStudy (kewlgeek555):

Yeah, let me solve the rest...

OpenStudy (kewlgeek555):

b equals 1.66...7

OpenStudy (anonymous):

right

OpenStudy (kewlgeek555):

I know what I did wrong! When I was solving c (before now) I solved it right to left, but I didn't remember that I have to do PEMDAS backwards! So c is -18 / 2 = -9 + 8 = -1. Yya!

OpenStudy (kewlgeek555):

Thank you both. Im sorry for wasting your times. I need to check my work more often.